what is the percent composition of oxygen in glucose, C6H12O6

1
Expert's answer
2021-03-08T05:55:54-0500

Answer: 53.3%

Molecular formula: C6H12O6.

Each glucose molecule contains:

6 oxygen atomsˆ

an Avogadro number of glucose molecules (1 mole of glucose) contains

6 × Avogadro number (= 6 moles) of O atoms.

1 mole of glucose has a mass (its gram formula weight or molar mass) which is the sum of the gram atomic weights of all of the constituent atoms.i.e. mass of 1 mole of glucose, C6H12O6 = (6 × 12.01 + 12 × 1.01 + 6 × 16.00) g = 180.18 g (using atomic weight data to 2 decimals)

1 mole of O atoms weighs 16.00 g and there are 6 moles of O atoms in 1 mole of glucose, so the mass of oxygen in 1 mole of glucose = 6 × 16.00 g = 96.00 g.


Percentage oxygen= 96.00/180.18 × 100 % = 53.3 %